Assume that Expansion Ltd is a diversified company that is considering an expansion project in a mining division. The company has a target debt-equity ratio of 1:2 and this ratio will not be affected by the new project. The company's manager has identified Dig-it-out Ltd as a company with the same business risk as the new project (equity beta of 1.5). Dig-it-out has a debt-equity ratio of 1:3. What is the beta estimate of Expansion Ltd?

A. 1.69
B. 1.12
C. 1.5
D. 1.75
Answer #1

Tax rate is not given so asusming it to be zero

unlevered beta=equity beta/(1+tax rate)*Debt/Equity)=1.5/(1+(1-0%)*1/3)=1.125

levered beta=unlevered beta*(1+(1-tax rate)*Debt/Equity)=1.125*(1+(1-0%)*1/2)=1.6875

Beta estimate
=1.69